OpenStudy (anonymous):

The reaction 2ClO2 + 2OH- -> ClO3- + ClO2 was studied with the following results: [ClO2](M) [OH-](M) Rate (M/s) 0.060 0.030 0.0248 0.020 0.030 0.00276 0.020 0.090 0.00828 a) Determine the rate law for the reaction b) Calculate the rate constant c) Calculate the rate when [ClO2] = 0.100 M and [OH-] = 0.050 M

OpenStudy (jadedry):

Rate of the reaction is given by the following equation: \[rate = k[A]^m[B]^n\] where k is the rate constant, m and n are some exponent. and A and B are the initial molar concentrations of the reactants involved. I see that temperature is not involved, so we can leave that out of our calculations and assume we are speaking of reactants at a constant temperature. For trials 1 and 2 we have: \[rate = k[0.060]^m[0.030]^n\] -1 \[rate = k[0.020]^m[0.030]^n\] -2 Divide the first rate by the second rate = 0.0248 /0.00276 = 8.98 = 9. therefore, the rate increased by a factor of 9 when ClO2 was increased. In dividing this way, you cancel out k and OH-^n since OH remained constant in both trials. = \[9 = 3^m\] therefore m = 2 for the trials 2 and 3 we have: \[rate = k[0.020]^m[0.030]^n\] -2 \[rate = k[0.020]^m[0.090]^n\] -3 Divide the second rate by the first rate = 0.00828/0.00276 = 3 Therefore the rate increased by a factor of 3 as OH- was increased. therefore we get: \[3=3^n \] hence n = 1. therefore our rate law is: \[Rate = k[ClO _{2}]^2[OH-]^1\] -- Answers to b) and c) coming in a second.

OpenStudy (jadedry):

In order to find the rate constant, we take our rate law: \[rate = k[ClO _{2}]^2[OH-]^1\] =\[\frac{ rate}{[ClO _{2}]^2[OH-]^1 }= k\] Take any values from the above table, I'll use trial one values: \[\frac{ 0.0248 }{ 0.060^2 *0.030^1 }\] \[= 229.62 = 2.3*10^2\] Therefore, our rate constant \[ = 2.3*10^2\] ---

OpenStudy (jadedry):

In order to find the rate when [ClO2] = 0.100 M and [OH-] = 0.050 M , simply plug in the values : \[rate = 2.29*10^2[0.100]^2[0.050]\] \[=0.1145\] = the rate. I hope this was helpful! c: